# Large Language Models Empowered Personalized Web Agents

### v1

@@ Abstract.
conference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYccs: Information systems Web applicationsccs: Information systems Personalization
@@ 1. Introduction
MiniWoB++ (Liu et al., 2018) Single-turn Mobile Web UI
RUSS (Xu et al., 2021) Multi-turn Web UI
META-GUI (Sun et al., 2022) Multi-turn Mobile apps
WebShop (Yao et al., 2022) Single-turn Shopping Web UI
Mind2Web (Deng et al., 2023) Single-turn Web UI
WebArena (Zhou et al., 2024) Single-turn Web UI
VWA (Koh et al., 2024) Single-turn Web UI
WebVoyager (He et al., 2024) Single-turn Web UI
WorkArena (Drouin et al., 2024) Multi-turn Web UI
WebLINX (Lù et al., 2024) Multi-turn Web UI
MT-Mind2Web (Deng et al., 2024) Multi-turn Web UI
MMInA (Zhang et al., 2024) Single-turn Web UI
Turking (Xu et al., 2024) Single-turn Web UI
ChatShop (Chen et al., 2024a) Multi-turn Web function
@@ 3.2. Benchmark Construction
It is challenging to gather a set of users to collect the real data on Web agent applications. Therefore, we chose to develop the benchmark using existing datasets of users’ Web behaviors. Specifically, we selected the Amazon Review (Hou et al., 2024) dataset as the foundation for our benchmark, as it provides a large-scale collection of users’ Web behaviors, including purchase and product rating across various categories of products. The overall pipeline for constructing PersonalWAB is illustrated in Figure A.3, which consists of three steps: personalized data construction, user instruction creation, and Web environment implementation.
@@ 3.3. Evaluation
    For review instructions, we assess the similarity between the agent’s generated review and the user’s actual review. We employ the sentence-transformer (Reimers and Gurevych, 2019) model to compute the cosine similarity, yielding a res acc between 0 and 1.
@@ 4.1. Statistic Analysis
$\bullet$ Instruction statistics. We examined the number and average token length of different instructions in Figure 4 (b). It is observed that the recommendation instructions have the smallest number of tokens because the recommendation is an exploratory task and doesn’t contain many user information requirements. The review instructions show slightly higher complexity compared to search and recommendation instructions, as they include many words for users to express their initial evaluations.
@@ 5.2. Function Parameter Optimization
Function Acc Res Acc Function Acc Res Acc Function Acc Res Acc Function Acc Res Acc
@@
Table 4. Performance comparison between our proposed method, PUMA, and baselines in multi-turn track. F. Acc represents function accuracy, R. Acc stands for result accuracy, and Avg. Steps indicate the average number of steps taken by the agent to complete each instruction.
Method Search Recommendation Review Overall
F. Acc R. Acc Avg. Steps F. Acc R. Acc Avg. Steps F. Acc R. Acc Avg. Steps F. Acc R. Acc Avg. Steps
No Memory 0.996 0.656 2.398 0.096 0.000 2.420 1.000 0.446 2.019 0.685 0.358 2.280
Random Memory 0.999 0.680 4.193 0.703 0.042 4.474 1.000 0.448 2.007 0.896 0.380 3.564
Last Memory 0.996 0.676 4.229 0.708 0.045 4.252 1.000 0.449 2.007 0.897 0.381 3.498
Relevant Memory 0.996 0.686 4.233 0.715 0.042 4.564 0.999 0.448 2.008 0.899 0.383 3.609
ReAct (Yao et al., 2023) 0.996 0.674 4.657 0.218 0.013 5.468 0.974 0.448 2.129 0.718 0.369 4.098
Reflection (Shinn et al., 2024) 1.000 0.686 5.406 0.281 0.014 6.145 0.976 0.449 2.145 0.741 0.373 4.579
RecMind (Wang et al., 2024) 0.997 0.642 6.728 0.347 0.026 6.003 0.997 0.451 2.107 0.771 0.364 4.938
InteRecAgent (Huang et al., 2023) 0.999 0.642 3.110 0.618 0.022 3.008 1.000 0.447 2.001 0.867 0.362 2.706
@@ 6. Experiments
We evaluate a range of baselines that employ different strategies for selecting and utilizing user history. These baselines are categorized into three groups: Memory Retrieval based Methods (No Memory, Random Memory, Last Memory, Relevant Memory), Enhanced Reasoning Methods (ReAct (Yao et al., 2023), Reflection (Shinn et al., 2024)), and Recommendation-Specific Memory Frameworks (Recmind (Wang et al., 2024), InteRecAgent (Huang et al., 2023)). The implementation details of baselines and our method are illustrated in § A.2 and § A.3.
@@ 6.1.2. Multi-turn Track
The multi-turn track results (Table 4) offer valuable insights into how different methods handle complex interactions. 1) First, baselines perform better in search and recommendation tasks compared to the single-turn track, benefiting from multiple attempts and user feedback, while review tasks show minimal improvement, as the agents typically follow a straightforward flow with limited feedback opportunities. 2) The memory retrieval baselines follow similar trends to the single-turn track, with relevant memory improving function accuracy and result accuracy, but at the cost of additional steps. 3) ReAct and Reflection perform worse than memory retrieval methods, requiring more steps and yielding lower accuracy. The complexity of these methods, which include reasoning and self-reflection, seems to hinder task efficiency and accuracy with extra input token length. 4) RecMind also requires a higher number of steps, as it performs additional function calls, but struggles with instruction identification. InteRecAgent uses fewer steps due to its streamlined memory, but this simplification results in lower result accuracy. 5) Our Task-specific Memory method performs strongly, particularly in search and recommendation tasks. By effectively extracting relevant information and filtering out redundant data, it enables more informed decisions with fewer steps. Although we did not evaluate the full PUMA approach due to model limitations in multi-turn settings, the results highlight the importance of task-specific memory in enhancing both efficiency and accuracy.
@@ 6.2. In-depth Analysis
We performed a comprehensive analysis of PUMA on ablation study, memory length, efficiency, action transitions, and multi-turn performance variation. The results for the first two are provided in § A.4 due to the limited scope.
@@
6.2.2. Analysis on action transitions.
6.2.3. Analysis of multi-turn performance variation.
Figure 8. Transitions of the agent’s actions in multi-turn interactions for search and recommendation instructions, respectively.
@@ A.1. Details of Profile Evaluation Experiments
We adopt the same setting with (Chen et al., 2024b), set the number of positive samples to 1 and 3, and negative samples to 4 and 7 in the user prediction and recommendation tasks, respectively. Experiments are conducted with gpt-4o-mini-2024-07-18, and results in Figure 5 show our profile exhibits significant improvements across both tasks, indicating that our profiles are more distinct and better aligned with user behaviors.
@@ A.2. Details of Baselines
Enhanced Reasoning Methods. We also tested frameworks designed to enhance the agent’s reasoning and decision-making capabilities. 1) ReAct: Proposed by (Yao et al., 2023), this framework instructs the language model to think before taking an action and generate “Thought: some reasoning Action: some JSON format action argument” to interact, enabling the model to deliberate over the information available and decide on the most appropriate action. 2) Reflection: Reflection (Shinn et al., 2024) builds on frameworks of ReAct by adding a self-evaluation phase, where the agent reviews and analyzes its previous actions and outcomes before proceeding. This process allows the agent to recognize mistakes, reassess decisions, and refine its strategy in subsequent interactions. We tested this baseline only in the multi-turn track, where the agent treats each user message as feedback for reflection and adjustment.
Recommendation-Specific Memory Frameworks. Recommendation tasks are inherently personalized, as they rely on a deep understanding of user preferences and behaviors. Given this, we include baselines that leverage memory mechanisms developed for recommendation agents, assessing their ability to enhance personalization in our context. 1) RecMind: An LLM-powered agent designed for general recommendation purposes (Wang et al., 2024), consists of two parts of memory, personalized memory, and world knowledge. Personalized Memory includes individualized user information, such as their reviews or ratings for a particular item. World Knowledge consists of item metadata information and real-time information accessed with a Web search function. In our setup, we retain the personalized memory containing user reviews and ratings, and we incorporate an additional function to enable RecMind to access detailed product information. 2) InteRecAgent: Proposed by (Huang et al., 2023), this framework uses LLMs as the core reasoning engine while utilizing recommender models as functions for interactive recommendations. Its memory structure includes a candidate bus (which stores current item candidates) and a user profile that captures three facets of user preferences: “like”, “dislike”, and “expect”. We adopt the user profile memory in our experiments and allow the agent to update this profile at the end of each task. As the user profile is synthesized by LLMs based on conversation history with the user, we evaluate this method only in the multi-turn setting, where ongoing dialogue allows for continuous adaptation of the user profile.
@@ A.3. Implementation Details
Baselines. We use gpt-4o-mini-2024-07-18 as the base language model across all baseline methods. For memory retrieval baselines, we set the memory length to 50 behaviors for the single-turn track and 20 behaviors for the multi-turn track, allowing additional input length for user messages and function results. For the Relevant Memory method, we calculate cosine similarity using the sentence-transformer (Reimers and Gurevych, 2019) to identify relevant behaviors. The ReAct baseline is combined with the Last Memory approach to ensure that reasoning processes have recent context, and we further extend this with a Reflection mechanism for multi-turn scenarios. For RecMind, the memory length is set to 400 behaviors, as it only contains user reviews and ratings, and we added an extra “get_product_details_by_asin” function for the agent to retrieve detailed product information. In the InteRecAgent setup, we first construct the memory using historical behaviors and the training dataset before evaluating performance on the test set.
@@ A.4. More Analysis
Ablation study. We conducted an ablation study (see Table 5) to assess the impact of PUMA’s key components. First, removing the memory leads to a significant drop in result accuracy across all tasks, highlighting the importance of memory in retaining relevant information for function parameter generation. Second, when memory is retained but the SFT phase is removed, result accuracy dramatically declines. This indicates that without fine-tuning, the model struggles to generate function parameters that align with user needs. Finally, removing the DPO phase results in a slight performance decrease, suggesting that DPO plays a crucial role in aligning the model with user preferences, and improving the quality of function parameters.
@@
Analysis on memory length. We evaluated the impact of different memory token lengths on our framework’s performance across tasks. The experiment measured both function accuracy and result accuracy with varying memory sizes. The results in Table 6 indicate that increasing memory length has minimal impact on function accuracy, with the model maintaining similar performance regardless of memory size. However, memory length has a significant effect on result accuracy, particularly in recommendation tasks. Shorter memory lengths reduce the number of stored products, limiting the model’s ability to select appropriate product IDs, which leads to a noticeable drop in recommendation accuracy. In contrast, search and review tasks are less sensitive to memory length changes, as the agent relies more on information from the instruction rather than the memory. This reduced dependence on memory in these tasks may also limit the model’s potential to improve performance further.
Appendix B Prompting Details

### v2

@@ Abstract.
journalyear: 2025copyright: ccconference: Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3696410.3714842isbn: 979-8-4007-1274-6/25/04ccs: Information systems Web applicationsccs: Information systems Personalization
@@ 1. Introduction
MiniWoB++ (Liu et al., 2018) Single-turn Mobile Web UI $\times$
RUSS (Xu et al., 2021) Multi-turn Web UI $\times$
META-GUI (Sun et al., 2022) Multi-turn Mobile apps $\times$
WebShop (Yao et al., 2022) Single-turn Shopping Web UI $\times$
Mind2Web (Deng et al., 2023) Single-turn Web UI $\times$
WebArena (Zhou et al., 2024) Single-turn Web UI $\times$
VWA (Koh et al., 2024) Single-turn Web UI $\times$
WebVoyager (He et al., 2024) Single-turn Web UI $\times$
WorkArena (Drouin et al., 2024) Multi-turn Web UI $\times$
WebLINX (Lù et al., 2024) Multi-turn Web UI $\times$
MT-Mind2Web (Deng et al., 2024) Multi-turn Web UI $\times$
MMInA (Zhang et al., 2024) Single-turn Web UI $\times$
Turking (Xu et al., 2024) Single-turn Web UI $\times$
ChatShop (Chen et al., 2024a) Multi-turn Web function $\times$
@@ 3.2. Benchmark Construction
It is challenging to gather a set of users to collect the real data on Web agent applications. Therefore, we chose to develop the benchmark using existing datasets of users’ Web behaviors. Specifically, we selected the Amazon Review (Hou et al., 2024) dataset as the foundation for our benchmark, as it provides a large-scale collection of users’ Web behaviors, including purchase and product rating across various categories of products. The overall pipeline for constructing PersonalWAB is illustrated in Figure 2, which consists of three steps: personalized data construction, user instruction creation, and Web environment implementation.
@@ 3.3. Evaluation
    For review instructions, we assess the similarity between the agent’s generated review and the user’s actual review. We employ the sentence-transformer (Reimers and Gurevych, 2019a) model to compute the cosine similarity, yielding a res acc between 0 and 1.
@@ 4.1. Statistic Analysis
$\bullet$ Instruction statistics. We examined the number and average token length of different instructions in Figure 4 (b). It is observed that the recommendation instructions have the smallest number of tokens because the recommendation is an exploratory task and doesn’t contain many user information requirements. The review instructions show slightly higher complexity than search and recommendation instructions, as they include many words for users to express their initial evaluations.
@@ 5.2. Function Parameter Optimization
Function Acc. Res Acc Function Acc. Res Acc Function Acc. Res Acc Function Acc. Res Acc
@@
Table 4. Performance comparison between our proposed method, PUMA, and baselines in multi-turn track. F. Acc. represents function accuracy, R. Acc. stands for result accuracy, and Avg. Steps indicate the average number of steps taken by the agent to complete each instruction.
Method (backbone) Search Recommendation Review Overall
F. Acc. R. Acc. Avg. Steps F. Acc. R. Acc. Avg. Steps F. Acc. R. Acc. Avg. Steps F. Acc. R. Acc. Avg. Steps
No Memory (gpt-4o) 0.996 0.656 2.398 0.096 0.000 2.420 1.000 0.446 2.019 0.685 0.358 2.280
Random Memory (gpt-4o) 0.999 0.680 4.193 0.703 0.042 4.474 1.000 0.448 2.007 0.896 0.380 3.564
Last Memory (gpt-4o) 0.996 0.676 4.229 0.708 0.045 4.252 1.000 0.449 2.007 0.897 0.381 3.498
Relevant Memory (gpt-4o) 0.996 0.686 4.233 0.715 0.042 4.564 0.999 0.448 2.008 0.899 0.383 3.609
ReAct (Yao et al., 2023) (gpt-4o) 0.996 0.674 4.657 0.218 0.013 5.468 0.974 0.448 2.129 0.718 0.369 4.098
Reflexion (Shinn et al., 2024) (gpt-4o) 1.000 0.686 5.406 0.281 0.014 6.145 0.976 0.449 2.145 0.741 0.373 4.579
RecMind (Wang et al., 2024) (gpt-4o) 0.997 0.642 6.728 0.347 0.026 6.003 0.997 0.451 2.107 0.771 0.364 4.938
InteRecAgent (Huang et al., 2023) (gpt-4o) 0.999 0.642 3.110 0.618 0.022 3.008 1.000 0.447 2.001 0.867 0.362 2.706
@@ 6. Experiments
We evaluate a range of baselines that employ different strategies for selecting and utilizing user history. These baselines are categorized into three groups: Memory Retrieval based Methods (No Memory, Random Memory, Last Memory, Relevant Memory), Enhanced Reasoning Methods (ReAct (Yao et al., 2023), Reflexion (Shinn et al., 2024)), and Recommendation-Specific Memory Frameworks (Recmind (Wang et al., 2024), InteRecAgent (Huang et al., 2023)). The implementation details of baselines and our method are illustrated in § A.2 and § A.3. All methods use the same prompt template, with differences only in the memory component. The detailed prompt design is provided in Figure 16 and Figure 17.
@@ 6.1.2. Multi-turn Track
The multi-turn track results (Table 4) offer valuable insights into how different methods handle complex interactions. 1) First, baselines perform better in search and recommendation tasks compared to the single-turn track, benefiting from multiple attempts and user feedback, while review tasks show minimal improvement, as the agents typically follow a straightforward flow with limited feedback opportunities. 2) The memory retrieval baselines follow similar trends to the single-turn track, with relevant memory improving function accuracy and result accuracy, but at the cost of additional steps. 3) ReAct and Reflexion perform worse than memory retrieval methods, requiring more steps and yielding lower accuracy. The complexity of these methods, which include reasoning and self-reflexion, seems to hinder task efficiency and accuracy with extra input token length. 4) RecMind also requires a higher number of steps, as it performs additional function calls, but struggles with instruction identification. InteRecAgent uses fewer steps due to its streamlined memory, but this simplification results in lower result accuracy. 5) Our Task-specific Memory method performs strongly, particularly in search and recommendation tasks. By extracting relevant information and filtering out redundant data, it enables more informed decisions with fewer steps. Although we did not evaluate the full PUMA approach due to model limitations in multi-turn settings, the results highlight the importance of task-specific memory in enhancing both efficiency and accuracy.
@@ 6.2. In-depth Analysis
We performed a comprehensive analysis of PUMA, including experiments on ablation study, memory length, efficiency, action transitions, multi-turn performance variation, function usage and outcome accuracy, search function implementation, and zero-shot and few-shot performance. Due to space limitations, we present the results for efficiency in this section, while the remaining analyses are provided in § A.4.
@@ 7. Conclusion and Future Work
Although this work makes significant strides in personalized Web agents, it is important to note some limitations. Ethical and privacy considerations, along with the scope of our study, warrant further discussion. These aspects are elaborated in § B, where we provide a detailed exploration of challenges associated with user data usage, potential biases in personalization, and the current focus on shopping scenarios. While our approach demonstrates strong adaptability, extending it to broader Web environments presents additional complexities that require further investigation.
Acknowledgements.
The work described in this paper was supported by the Research Grants Council of Hong Kong (PolyU/15209724, PolyU/15207821, PolyU/15207122, PolyU/15213323) and PolyU internal grants (BDWP).
@@ A.1. Details of Profile Evaluation Experiments
We adopt the same setting with (Chen et al., 2024b), set the number of positive samples to 1 and 3, and negative samples to 4 and 7 in the user prediction and recommendation tasks, respectively. To ensure a fair evaluation, both experiments are conducted in separate, isolated sessions. Experiments are conducted with gpt-4o-mini-2024-07-18, and results in Figure 5 show our profile exhibits significant improvements across both tasks, indicating that our profiles are more distinct and better aligned with user behaviors.
@@ A.2. Details of Baselines
Enhanced Reasoning Methods. We also tested frameworks designed to enhance the agent’s reasoning and decision-making capabilities. 1) ReAct: Proposed by (Yao et al., 2023), this framework instructs the language model to think before taking an action and generate “Thought: some reasoning Action: some JSON format action argument” to interact, enabling the model to deliberate over the information available and decide on the most appropriate action. 2) Reflexion: Reflexion (Shinn et al., 2024) builds on frameworks of ReAct by adding a self-evaluation phase, where the agent reviews and analyzes its previous actions and outcomes before proceeding. This process allows the agent to recognize mistakes, reassess decisions, and refine its strategy in subsequent interactions. We tested this baseline only in the multi-turn track, where the agent treats each user message as feedback for Reflexion and adjustment.
Recommendation-Specific Memory Frameworks. Recommendation tasks are inherently personalized, as they rely on a deep understanding of user preferences and behaviors (Kang and McAuley, 2018; Huang et al., 2023; Wang et al., 2024; Gao et al., 2023; Li et al., 2024). Given this, we include baselines that leverage memory mechanisms developed for recommendation agents, assessing their ability to enhance personalization in our context. 1) RecMind: An LLM-powered agent designed for general recommendation purposes (Wang et al., 2024), consists of two parts of memory, personalized memory, and world knowledge. Personalized Memory includes individualized user information, such as their reviews or ratings for a particular item. World Knowledge consists of item metadata information and real-time information accessed with a Web search function. In our setup, we retain the personalized memory containing user reviews and ratings, and we incorporate an additional function to enable RecMind to access detailed product information. 2) InteRecAgent: Proposed by (Huang et al., 2023), this framework uses LLMs as the core reasoning engine while utilizing recommender models as functions for interactive recommendations. Its memory structure includes a candidate bus (which stores current item candidates) and a user profile that captures three facets of user preferences: “like”, “dislike”, and “expect”. We adopt the user profile memory in our experiments and allow the agent to update this profile at the end of each task. As the user profile is synthesized by LLMs based on conversation history with the user, we evaluate this method only in the multi-turn setting, where ongoing dialogue allows for continuous adaptation of the user profile.
@@ A.3. Implementation Details
Baselines. We use gpt-4o-mini-2024-07-18 as the base language model across all baseline methods. For memory retrieval baselines, we set the memory length to 50 behaviors for the single-turn track and 20 behaviors for the multi-turn track, allowing additional input length for user messages and function results. For the Relevant Memory method, we calculate cosine similarity using the sentence-transformer (Reimers and Gurevych, 2019a) to identify relevant behaviors. The ReAct baseline is combined with the Last Memory approach to ensure that reasoning processes have recent context, and we further extend this with a Reflexion mechanism for multi-turn scenarios. For RecMind, the memory length is set to 400 behaviors, as it only contains user reviews and ratings, and we added an extra “get_product_details_by_asin” function for the agent to retrieve detailed product information. In the InteRecAgent setup, we first construct the memory using historical behaviors and the training dataset before evaluating performance on the test set.
@@ A.4. More Analysis
We conducted an ablation study (see Table 5) to assess the impact of PUMA’s key components. First, removing the memory leads to a significant drop in result accuracy across all tasks, highlighting the importance of memory in retaining relevant information for function parameter generation. Second, when memory is retained but the SFT phase is removed, result accuracy dramatically declines. This indicates that without fine-tuning, the model struggles to generate function parameters that align with user needs. Finally, removing the DPO phase results in a slight performance decrease, suggesting that DPO plays a crucial role in aligning the model with user preferences, and improving the quality of function parameters.
A.4.2. Analysis on memory length.
@@
We evaluated the impact of different memory token lengths on our framework’s performance across tasks. The experiment measured both function accuracy and result accuracy with varying memory sizes. The results in Table 6 indicate that increasing memory length has minimal impact on function accuracy, with the model maintaining similar performance regardless of memory size. However, memory length has a significant effect on result accuracy, particularly in recommendation tasks. Shorter memory lengths reduce the number of stored products, limiting the model’s ability to select appropriate product IDs, which leads to a noticeable drop in recommendation accuracy. In contrast, search and review tasks are less sensitive to memory length changes, as the agent relies more on information from the instruction rather than the memory. This reduced dependence on memory in these tasks may also limit the model’s potential to improve performance further.
A.4.3. Analysis on action transitions.
Figure 8. Transitions of the agent’s actions in multi-turn search and recommendation tasks. Each color represents a specific function. The horizontal axis shows interaction steps, while the width of each color band indicates the proportion of the agent’s focus on that action. The flow between steps illustrates how the agent adapts its strategy over steps.
A.4.4. Analysis of multi-turn performance variation.
Table 7. Single-turn performance comparison of different methods in terms of function accuracy (F. Acc.), result accuracy (R. Acc.), and outcome accuracy (O. Acc.) in search and recommendation.
Method Search Recommendation
F. Acc. R. Acc. O. Acc. F. Acc. R. Acc. O. Acc.
No Memory 1.000 0.647 0.647 0.092 0.000 0.155
Random Memory 0.974 0.640 0.642 0.296 0.018 0.159
Last Memory 0.937 0.626 0.632 0.432 0.028 0.161
Relevant Memory 0.928 0.622 0.631 0.492 0.030 0.159
ReAct (Yao et al., 2023) 0.903 0.605 0.628 0.560 0.027 0.160
RecMind (Wang et al., 2024) 0.981 0.645 0.647 0.226 0.017 0.152
PUMA 1.000 0.649 0.649 0.939 0.048 0.164
Table 8. Comparison of search result accuracy using BM25 and Dense retrieval methods in single-turn track.
Method Search (Result Accuracy)
BM25 Dense Retrieval
No Memory 0.647 0.502
Random Memory 0.640 0.504
Last Memory 0.626 0.498
Relevant Memory 0.622 0.499
ReAct (Yao et al., 2023) 0.605 0.496
RecMind (Wang et al., 2024) 0.645 0.498
PUMA 0.649 0.506
Table 9. Performance comparison in zero-shot and few-shot scenarios in single-turn track.
Method Search Recommendation Review Overall
Function Acc. Result Acc. Function Acc. Result Acc. Function Acc. Result Acc. Function Acc. Result Acc.
No Memory 1.000 0.684 0.050 0.000 1.000 0.388 0.625 0.328
Random Memory 0.974 0.684 0.301 0.060 0.996 0.391 0.715 0.352
Last Memory 1.000 0.683 0.314 0.058 1.000 0.396 0.730 0.353
Relevant Memory 0.928 0.675 0.405 0.078 1.000 0.397 0.743 0.358
ReAct (Yao et al., 2023) 0.945 0.675 0.475 0.080 0.996 0.393 0.774 0.358
RecMind (Wang et al., 2024) 0.973 0.680 0.320 0.063 0.996 0.394 0.722 0.354
PUMA 1.000 0.686 0.892 0.090 1.000 0.396 0.958 0.366
A.4.5. Analysis on function usage and outcome accuracy.
In real-world applications, users care about whether the system retrieves relevant results rather than the specific function it employs. While search and recommendation serve different roles, their distinction may not always be meaningful from a user’s perspective. For real users, the ultimate priority lies in the relevance of the returned items rather than the specific function used by the agent.
To better align with this user-centric goal, we introduce Outcome Accuracy (O. Acc.), a metric that evaluates the correctness of the returned results independent of the function used for search and recommendation tasks. As shown in Table 7, function accuracy (F. Acc.) and result accuracy (R. Acc.) vary significantly between search and recommendation tasks. However, Outcome Accuracy provides a more balanced perspective on the effectiveness of different methods. Our results show that PUMA achieves the highest Outcome Accuracy across both search and recommendation tasks, demonstrating its ability to deliver relevant results without being constrained by function selection. By incorporating this metric, we offer a more comprehensive evaluation that prioritizes relevance over strict function adherence, reflecting real-world user needs.
A.4.6. Analysis on search function implementation.
In our benchmark, the implementations of both the search and recommendation functions are flexible and not restricted to a specific method. Our focus is not on optimizing these functions but on evaluating the agent’s ability to effectively utilize the provided function in a personalized setting.
We conducted an alternative retrieval experiment in the single-turn track, replacing BM25 with a dense retrieval model based on Sentence-BERT (Reimers and Gurevych, 2019b). The results in Table 8 indicated that while dense retrieval captures richer semantic representations, it also introduces noise by embedding extensive product details, leading to a slight degradation in result accuracy across all baselines. And different retrieval methods may impact performance, but the overall trends remain consistent. Despite these variations, our PUMA framework consistently outperformed all baselines, regardless of the retrieval model used. Given the modular nature of our benchmark, future work could explore alternative retrieval models and different recommendation strategies, further analyzing how function-level improvements lead to final performance.
A.4.7. Analysis on zero-shot and few-shot performance.
Since our user sampling is fully random, some users naturally have limited historical data, leading to zero-shot and few-shot scenarios. To evaluate performance in such cases, we analyzed 139 users with fewer than 10 historical records (16.2% of the test set) in the single-turn track. Results in Table 9 show task-dependent effects: Search performance remains stable or improves slightly due to reduced irrelevant information, while Recommendation performance also improves, as limited memory simplifies retrieval. However, Review tasks decline, as the lack of past reviews hinders personalized responses. Despite these variations, PUMA consistently outperforms all baselines, achieving the highest accuracy across tasks, particularly excelling in Recommendation scenarios. These results demonstrate that PUMA remains effective even in zero-shot and few-shot conditions, highlighting its adaptability when user history is sparse.
Appendix B Discussion
B.1. Ethical and Privacy Considerations
The integration of personalized data in Web agents plays a crucial role in improving task efficiency and relevance. By leveraging user history, preferences, and past interactions, personalized Web agents can better align with individual needs and enhance task completion. However, the use of such data raises important ethical and privacy concerns that must be carefully considered.
One primary ethical concern is fairness in personalization. If personalization is not properly managed, it may introduce or amplify biases in decision-making processes. For example, recommendation algorithms might reinforce popularity bias (Abdollahpouri et al., 2019), leading to a concentration of exposure on frequently suggested items while limiting diversity. Additionally, personalization can inadvertently discriminate against certain user groups if historical biases exist in the training data (Biega et al., 2018). To mitigate these risks, fairness-aware personalization techniques and diversity-promoting strategies should be considered when designing personalized Web agents.
From a privacy perspective, handling user history introduces risks related to data security. Users may inadvertently disclose sensitive information, such as browsing behaviors, or purchase histories, which could be exploited if not properly secured. To ensure privacy protection, future work about personalized Web agents should incorporate privacy-preserving techniques, reducing the risk of data leakage while maintaining personalization benefits.
B.2. Scope of This Work
We chose the shopping domain as a representative and practically significant application to demonstrate the capabilities of personalized Web agents. Shopping platforms provide structured user behavior data and diverse interactions, such as search, recommendation, and product reviews, making them ideal testbeds for evaluating personalization strategies. Additionally, personalization in shopping directly impacts user experience and decision-making, highlighting the effectiveness of adaptive Web agents.
While our benchmark is designed on shopping-related Web functions, the framework introduced in this work is generalizable and can be extended to a broader range of Web applications, such as news recommendation, and social media content curation. These domains, like shopping, rely on user behavior modeling to enhance relevance and improve interaction quality.
Appendix C Prompting Details